OpenStudy (anonymous):

A singing competition is auditioning 35 contestants each hour for the next h hours. Which equation can be used to find the total number of contestants, C, that audition in h hours?

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

Are there answer choices?

OpenStudy (anonymous):

OK, hold on one second.

OpenStudy (anonymous):

35h = C (1.) 35 + h = C (2.) 35 -- = C (3.) h h -- = C (4.) 35

OpenStudy (anonymous):

@SACAPUNTAS can you help me and shamil98

OpenStudy (anonymous):

Do you guy's know?

OpenStudy (anonymous):

There are 35 per hour for \(h\) hours. So you just multiply them. It's the same as if you were driving 55 mph for h hours and you wanted to know how far you went.

OpenStudy (anonymous):

Can you pick one of my choice.

OpenStudy (anonymous):

nevermind